# Implicit geometric regularization in flow matching via density weighted Stein operators

### v1

@@ 3.4 Geometric Foundation: The $\gamma$-Stein Viewpoint
The $\gamma$-Stein Metric. From the perspective of Information Geometry (eguchi2009information), the $\gamma$-divergence generates a Riemannian metric structure on the space of probability densities. Following standard conventions in information geometry, we hereinafter denote the parametric family of model densities by $q_{\theta}$. The weighting $w_{\gamma}\approx q_{\theta}^{\gamma}$ in our objective (Eq. 1) naturally corresponds to the $\gamma$-weighted Fisher information metric $g^{(\gamma)}$ (Fujisawa2008; Matsuzoe2017):
@@
Therefore, minimizing $\mathcal{L}_{\gamma}$ over admissible flows $\bigl(q_{\theta(t)},u_{t}\bigr)_{t\in[0,1]}$ can be interpreted as searching for geodesic curves on the statistical manifold endowed with the $\gamma$-Stein metric and connection. The case $\gamma=0$ reduces to the flat $L^{2}$ geometry, where the tangent space coincides with $L^{2}(q_{\theta(t)})$ and $\Pi_{\theta(t),0}$ is the identity, so that geodesics correspond to straight lines and the velocity field must be defined everywhere. For $\gamma>0$, the metric is weighted by $q_{\theta(t)}^{1+\gamma}$, so that directions supported in low-density regions have negligible norm. As a consequence, the optimal velocity field produced by $\gamma$-FM concentrates on the high-density manifold of the data, not because it “ignores” data, but because the intrinsic geometry itself is dominated by the manifold structure through the $\gamma$-Stein projection.
@@ 3.5 Implicit Geometric Regularization
$\mathcal{L}_{\text{roughness}}(t)\coloneqq\int_{\mathbb{R}^{d}}q_{\theta(t)}(x)w_{\gamma}(x,t)\|\nabla_{x}v_{\theta}(x,t)\|_{F}^{2}dx.$ (11)
@@
where $p_{\mathrm{ambient}}$ represents the distribution of the void regions (e.g., uniform noise in the bounding box). This reduction in the local Lipschitz constant implies a less stiff ODE, permitting adaptive solvers to take larger integration steps. This mechanism directly accounts for the improved NFE (Number of Function Evaluations) reported in our experiments.
@@
To connect the empirical reduction in (11) with a more geometric picture, it is convenient to introduce the weighted Dirichlet form associated with the escort weight. Such variational limits of discrete regularizers on data manifolds have been rigorously studied in the context of Optimal Transport by hamm2025manifold. Fix a time $t$ and write $p_{t}$ for the marginal density of $x_{t}$. Using the weighting factor $w_{\gamma}(x,t)\propto q_{\theta(t)}(x)^{\gamma}$, we introduce the escort measure
@@
which is the population analogue of a $\gamma$-FM regression loss with an explicit weighted Sobolev penalty. The next proposition studies an idealized population objective where we add an explicit weighted Sobolev penalty to the $\gamma$-FM loss. Although Algorithm 1 does not explicitly add a Tikhonov penalty, the analysis above identifies a roughness functional naturally associated with the $\gamma$-weighted objective. In practice, this provides a mechanistic explanation for the observed smoothness improvements, which may be further amplified by the implicit bias of SGD (see, e.g., Rahaman et al., 2019; Xu et al., 2020; Jin and Montúfar, 2023). The Dirichlet-regularized objective in (15) should therefore be viewed as a tractable surrogate that makes the geometric effect of the escort weighting explicit in the eigenbasis of the weighted Laplacian. Let $f_{\gamma,t}^{\ast}$ denote the unique minimizer of $\mathcal{J}_{\gamma,t}$.
@@
In the case of $\gamma=0$, the measure $\mu_{0,t}$ reduces to the standard density $q_{\theta(t)}$, and Eq. (18) recovers the standard spectral filtering result known in manifold regularization (belkin2005manifold). The significance of Proposition 3.4 lies in the dependence on $\gamma$: as discussed in the proof, increasing $\gamma$ effectively rescales the eigenvalues $\mu_{k}^{(\gamma,t)}$, thereby intensifying the shrinkage effect on high-frequency modes compared to the standard case. We suggest a close relation to the Witten Laplacian and the Bakry–Émery curvature in the proof. Assume that the marginal density $q_{\theta(t)}$ admits a smooth potential $U_{t}$ such that $q_{\theta(t)}(x)=\exp(-U_{t}(x))$. Then the generator $L_{\gamma,t}$ in (13) can be rewritten as
@@
$\mathcal{L}_{\text{roughness}}=\int_{\mathbb{R}^{d}}w_{\gamma}(x,t)\,\|\nabla_{x}v_{\theta}(x,t)\|_{F}^{2}\,\mathrm{d}x$ (20)
@@ 3.6 Theoretical Selection of $\gamma$ via Geometric Selection Criterion
$\mathrm{GSC}(\gamma)={\mathrm{MMD}^{2}(p_{\mathrm{data}},p_{\theta_{\gamma}})}+\lambda{\mathcal{R}_{\mathrm{smooth}}(v_{\theta_{\gamma}})},$ (21)
where $\lambda>0$ is a trade-off parameter (we set $\lambda=1$). Here, the squared Maximum Mean Discrepancy (MMD) serves as the bias proxy $L(\gamma)$, measuring generation quality. The roughness functional $\mathcal{R}_{\mathrm{roughness}}(v_{\theta_{\gamma}})$ derived in (20) serves as the stability penalty. Minimizing the GSC identifies a $\gamma$ that achieves the optimal trade-off between faithful data reconstruction and geometric regularity.
@@ 4.1 Synthetic Verification: Implicit Regularization
Experimental Setup. To simulate this regime, we construct a 2-dimensional ring manifold embedded in a high-dimensional space $\mathbb{R}^{20}$. The first two dimensions contain the data structure (a noisy circle), while the remaining 18 dimensions consist of low-magnitude Gaussian noise, representing the ambient space. We train both Standard FM ($\gamma=0$) and $\gamma$-FM ($\gamma=1$) on this dataset and analyze the learned vector fields by taking a 2D slice of the 20D space.
Results: Void Rejection and Energy Efficiency. Figure 2 visualizes the velocity norm $\|v_{\theta}(x)\|_{2}$ of the learned fields.
  • Standard FM ($\gamma=0$): As shown in the top row, the model learns a vector field that remains active with high magnitude even in the center of the ring (the void), where no data exists. This confirms that unweighted regression wastes model capacity by attempting to fit target signals in irrelevant regions, leading to a "hallucination" of flow in the empty space.
  • $\gamma$-FM ($\gamma=1$): In contrast, the bottom row shows that $\gamma$-FM effectively suppresses the vector field in the void. The velocity norm drops to near zero in the center (visualized as the dark region), and the flow is strictly confined to the vicinity of the data manifold.
This result empirically supports the finite-propagation intuition derived in Proposition 3.2. By ignoring the void, $\gamma$-FM implicitly regularizes the geometry, ensuring that the ODE solver does not waste evaluations on non-existent paths.
Figure 2: Visualization of Implicit Geometric Regularization in High Dimensions. We trained flow matching models on a 2D ring embedded in a 20-dimensional space ($D=20$). The plots show the 2D slice of the learned vector fields and their velocity norms. (Top) Standard FM: The vector field is active with high energy even in the data-free void (center), indicating inefficient global regression. (Bottom) $\gamma$-FM ($\gamma=1$): The density-weighted objective successfully suppresses the flow in the void (dark region in the rightmost heatmap), concentrating the vector field solely on the data manifold. This confirms the theoretical prediction of void rejection.
@@ 4.2 Latent-flow modelling of CIFAR-10
Our experimental setting follows the general latent-flow paradigm of Lipman2023 in the sense that we train a flow in the latent space of a pre-trained autoencoder rather than directly in pixel space. Concretely, we first train an autoencoder on CIFAR-10 and then freeze the encoder and decoder; all flow-matching experiments are carried out in the resulting latent space.
@@
In contrast to Lipman2023, who employ the standard (unweighted) flow-matching objective and focus on high-resolution image synthesis, we keep the latent architecture fixed and vary only the regression geometry in latent space via the $\gamma$-weighted loss (1). Thus, differences in Maximum Mean Discrepancy (MMD), smoothness, and NFE observed in Table 1 can be attributed to the effect of the $\gamma$-weighting rather than to changes in the autoencoder or flow architecture.
We evaluate $\gamma$-FM on the CIFAR-10 dataset using a latent flow model. An autoencoder compresses the images into a lower-dimensional latent space, and the flow is trained to model the distribution of these latents. We focus on two metrics:
  • RBF-MMD2 (lower is better): Measures the discrepancy between generated samples and real data in latent space.
  • Smoothness (lower is better): Measures the average squared Frobenius norm of the Jacobian of $v_{\theta}$, i.e., $\mathcal{R}_{\mathrm{smooth}}(v_{\theta})$. We report an ambient smoothness proxy under $\mathcal{N}(0,I)$ for stability comparison.
Table 1 summarizes the performance of $\gamma$-FM with varying $\gamma$. We evaluate the generation quality using RBF-MMD2 computed against the inlier and outlier evaluation sets. Additionally, we report the Smoothness of the learned vector field, defined as $\mathbb{E}_{x\sim\mathcal{N}(0,I)}[\|\nabla v(x)\|_{F}^{2}]$, where lower values indicate a smoother and more stable flow.
The results show that $\gamma=1.0$ achieves the best performance across all metrics. It yields the lowest Inlier MMD ($0.0126$), significantly outperforming the baseline ($\gamma=0.0$, $0.0481$). Furthermore, $\gamma=1.0$ achieves the lowest Smoothness score ($14.46$). This indicates that our weighting scheme not only filters out outliers but also regularizes the vector field, leading to a smoother flow that is easier to simulate. In contrast, excessively large $\gamma$ (e.g., $\gamma=4.0$) degrades both generation quality and smoothness, likely due to over-concentration of the probability density.
Table 1: Quantitative results on CIFAR-10 latent flow matching. We report RBF-MMD2 (lower is better) and Vector Field Smoothness (lower is better).
$\gamma$ Inlier MMD Outlier MMD Smoothness
0.0 (Baseline) 0.0481 0.0875 22.42
0.2 0.0490 0.0903 28.72
0.5 0.0299 0.0675 26.72
1.0 0.0126 0.0406 14.46
2.0 0.0466 0.0874 24.21
4.0 0.0485 0.0891 22.82
Adaptive Selection via GSC.
To select the optimal $\gamma$ without relying on visual inspection, we utilize the Geometric Selection Criterion (GSC) derived in Section 3.6. As shown in Figure 3(b), we evaluate the GSC (Eq. 21) across a grid of $\gamma$ values. The curve confirms that $\gamma=1.0$ effectively minimizes the combined objective, balancing the reduction in MMD (Bias) with the improvement in vector field smoothness (Variance). This theoretically grounded selection aligns with our quantitative results in Table 1, where $\gamma=1.0$ achieves the best generation quality.

### v2

@@ 3.4 Geometric Foundation: The $\gamma$-Stein Viewpoint
The $\gamma$-Stein Metric. From the perspective of Information Geometry (Eguchi, 2009), the $\gamma$-divergence generates a Riemannian metric structure on the space of probability densities. The $\gamma$-Stein operator for robust inference with unnormalized models was studied in Eguchi (2025). Following standard conventions in information geometry, we hereinafter denote the parametric family of model densities by $q_{\theta}$. The weighting $w_{\gamma}\approx q_{\theta}^{\gamma}$ in our objective (Eq. 2) naturally corresponds to the $\gamma$-weighted Fisher information metric $g^{(\gamma)}$ (Fujisawa and Eguchi, 2008; Matsuzoe, 2017):
@@
In this idealized geometry, minimizing $\mathcal{L}_{\gamma}$ over admissible flows $\bigl(q_{\theta(t)},u_{t}\bigr)_{t\in[0,1]}$ may be interpreted as searching for geodesic-like curves on the statistical manifold endowed with the $\gamma$-Stein metric and connection. The case $\gamma=0$ reduces to the flat $L^{2}$ geometry, where the tangent space coincides with $L^{2}(q_{\theta(t)})$ and $\Pi_{\theta(t),0}$ is the identity, so that geodesics correspond to straight lines and the velocity field must be defined everywhere. For $\gamma>0$, the metric is weighted by $q_{\theta(t)}^{1+\gamma}$, so that directions supported in low-density regions have negligible norm. This does not prove that the implemented algorithm exactly follows geodesics of the escort geometry; rather, it motivates why the learned velocity field tends to concentrate on the high-density manifold through the $\gamma$-Stein projection.
@@ 3.5 Implicit Geometric Regularization
$\mathcal{R}_{\text{roughness}}(t)\coloneqq\int_{\mathbb{R}^{d}}q_{\theta(t)}(x)w_{\gamma}(x,t)\|\nabla_{x}v_{\theta}(x,t)\|_{F}^{2}dx.$ (12)
@@
where $p_{\mathrm{ambient}}$ represents the distribution of the void regions (e.g., uniform noise in the bounding box). This reduction in the local Lipschitz constant implies a less stiff ODE, permitting adaptive solvers to take larger integration steps. This mechanism is consistent with the improved NFE behavior observed in our experiments, although we do not claim a universal causal guarantee beyond the tested setting.
@@
To connect the empirical reduction in (12) with a more geometric picture, it is convenient to introduce the weighted Dirichlet form associated with the escort weight. Such variational limits of discrete regularizers on data manifolds have been rigorously studied in the context of Optimal Transport by Hamm et al. (2025). Fix a time $t$ and write $p_{t}$ for the marginal density of $x_{t}$. Using the weighting factor $w_{\gamma}(x,t)\propto q_{\theta(t)}(x)^{\gamma}$, we introduce the escort measure
@@
which is the population analogue of a $\gamma$-FM regression loss with an explicit weighted Sobolev penalty. The next proposition studies an idealized population objective where we add an explicit weighted Sobolev penalty to the $\gamma$-FM loss. Although Algorithm 1 does not explicitly add a Tikhonov penalty, the analysis above identifies a roughness functional naturally associated with the $\gamma$-weighted objective. In practice, this provides a mechanistic explanation for the observed smoothness improvements, which may be further amplified by the implicit bias of SGD (see, e.g., Rahaman et al. (2019); Xu et al. (2020); Jin and Montúfar (2023)). The Dirichlet-regularized objective in (16) should therefore be viewed as a tractable surrogate that makes the geometric effect of the escort weighting explicit in the eigenbasis of the weighted Laplacian. Let $f_{\gamma,t}^{\ast}$ denote the unique minimizer of $\mathcal{J}_{\gamma,t}$.
@@
In the case of $\gamma=0$, the measure $\mu_{0,t}$ reduces to the standard density $q_{\theta(t)}$, and Eq. (19) recovers the standard spectral filtering result known in manifold regularization (Belkin et al., 2005). The significance of Proposition 3.4 lies in the dependence on $\gamma$: as discussed in the proof, increasing $\gamma$ effectively rescales the eigenvalues $\mu_{k}^{(\gamma,t)}$, thereby intensifying the shrinkage effect on high-frequency modes compared to the standard case. We suggest a close relation to the Witten Laplacian and the Bakry–Émery curvature in the proof. Assume that the marginal density $q_{\theta(t)}$ admits a smooth potential $U_{t}$ such that $q_{\theta(t)}(x)=\exp(-U_{t}(x))$. Then the generator $L_{\gamma,t}$ in (14) can be rewritten as
@@
$\mathcal{R}_{\text{roughness}}=\int_{\mathbb{R}^{d}}q_{\theta(t)}(x)w_{\gamma}(x,t)\,\|\nabla_{x}v_{\theta}(x,t)\|_{F}^{2}\,\mathrm{d}x$ (21)
@@ 3.6 Theoretical Selection of $\gamma$ via Geometric Selection Criterion
$\mathrm{GSC}(\gamma)={\mathrm{MMD}^{2}(p_{\mathrm{data}},p_{\theta_{\gamma}})}+\lambda{\mathcal{R}_{\mathrm{roughness}}(v_{\theta_{\gamma}})},$ (22)
where $\lambda>0$ is a trade-off parameter (we set $\lambda=1$). Here, the squared Maximum Mean Discrepancy (MMD) serves as the bias proxy, measuring generation quality. The roughness functional $\mathcal{R}_{\mathrm{roughness}}(v_{\theta_{\gamma}})$ derived in (21) serves as the stability penalty. Minimizing the GSC may help identify a $\gamma$ that balances faithful data reconstruction and geometric regularity.
@@ 4.1 Synthetic Verification: Implicit Regularization
Experimental Setup. To make the weighting mechanism visually transparent, we construct a ring-shaped target distribution in a high-dimensional ambient space $\mathbb{R}^{20}$. The first two coordinates contain the ring structure, while the remaining coordinates consist of low-level Gaussian noise. We then train both standard FM $(\gamma=0)$ and $\gamma$-FM $(\gamma=1)$, and visualize a two-dimensional slice of the learned fields. For this toy experiment, we use the same practical shared-time ordering surrogate as in the main experiments, and inspect a late interpolation slice at $t=0.85$ so that the ring-support region, the center void, and the outside-support region are visually separated.
Results: Where the weight acts. Figure 2 is designed to separate three distinct objects: the location of path mass, the location where the practical weighting proxy acts, and the geometry of the resulting learned field. Panel (a) shows that at the displayed late time the interpolation-path density is concentrated near the ring. Panel (b) shows the corresponding rank-based monotone weighting proxy, which places larger emphasis near that support region and downweights both the center void and the far exterior. Panels (c) and (d) then compare the learned speed maps under FM and $\gamma$-FM, while panel (e) summarizes the same comparison through radial profiles.
The main qualitative message is not that the practical surrogate is a literal density estimator, but that it provides a useful within-batch ordering signal. In this toy example, that ordering concentrates learning pressure near the ring-support region and weakens it in low-density regions. As a result, the learned field under $\gamma$-FM is visually smoother in the center void and the outside-support region than the field learned by standard FM. This revised visualization therefore clarifies where the weighting acts and how it changes the learned geometry, which was less transparent in the previous version of the figure.
Figure 2: Where the shared-time weighting proxy acts in the ring toy experiment. Panel (a) shows a two-dimensional slice of the interpolation-path density at a late shared time $t=0.85$, where the path mass is concentrated near the ring support. Panel (b) shows the corresponding practical rank-based weighting proxy $w_{\gamma}(x_{t})$, which places larger emphasis near the ring-support region and downweights both the center void and the outside-support region. Panels (c) and (d) compare the learned speed maps of standard FM and $\gamma$-FM, respectively. Panel (e) shows radial profiles of the path density, the weighting proxy, and the learned speeds. The figure is intended to separate where the path places mass, where the weighting proxy acts, and how the learned field changes as a result.
@@ 4.2 Latent-flow modelling of CIFAR-10
Our experimental setting follows the general latent-flow paradigm of Lipman et al. (2023) in the sense that we train a flow in the latent space of a pre-trained autoencoder rather than directly in pixel space. Concretely, we first train an autoencoder on CIFAR-10 and then freeze the encoder and decoder; all flow-matching experiments are carried out in the resulting latent space.
@@
In contrast to Lipman et al. (2023), who employ the standard (unweighted) flow-matching objective and focus on high-resolution image synthesis, we keep the latent architecture fixed and vary only the regression geometry in latent space via the $\gamma$-weighted loss (2). Thus, the comparisons in Table 1 and the accompanying discussion are designed so that differences in MMD, smoothness, NFE, and decoded-image metrics can be attributed to the effect of the $\gamma$-weighting rather than to changes in the autoencoder or flow architecture.
4.3 Evaluation on latent CIFAR-10
To address whether the observed gains of $\gamma$-FM arise specifically from density-weighted geometry rather than from generic smoothing alone, we expand the latent CIFAR-10 evaluation in four directions. Throughout this subsection, we work under a robust-target latent protocol rather than a purely clean one: the flow model is trained on an inlier–outlier mixture in latent space, while decoded-image metrics are evaluated against a held-out inlier-class test set. This setting allows us to study whether density weighting organizes the regression geometry toward the dominant data manifold, instead of merely acting as a generic smoother. As a complementary sanity check, we also repeated the same latent CIFAR-10 pipeline under a clean-target protocol, in which the flow model is trained on inlier latents only. That rerun is useful for separating two questions: whether the practical weighting mechanism remains operational on clean data, and whether it yields a measurable gain when the regression geometry is not distorted by contaminated or off-support targets. Empirically, the weighting remained active in the clean rerun, but the performance gain largely disappeared, with shared-time FM already competitive in FID, Recall, and Coverage. We therefore use the robust-target setting as the main empirical regime for evaluating the practical advantage of the proposed method.
First, we build the training protocol so that the dynamic density proxy is computed from a shared-time minibatch, as described in Section 2.3 and Algorithm 1. This ensures that the $k$-NN statistic is computed within a particle cloud sampled from a single marginal $p_{t}$ induced by the robust-target training distribution.
Second, we add an explicit regularization baseline:
$\mathrm{FM+Jac},$
which augments the standard flow-matching loss with a Jacobian penalty,
$\mathcal{L}_{\mathrm{FM+Jac}}(\theta)=\mathcal{L}_{\mathrm{FM}}(\theta)+\lambda_{J}\,\mathbb{E}_{t,x,\varepsilon}\bigl[\|J_{x}v_{\theta}(x,t)^{\top}\varepsilon\|^{2}\bigr],$
where $\varepsilon\sim\mathcal{N}(0,I)$ (or a Rademacher probe) and the Jacobian norm is estimated by a Hutchinson-type trace estimator. This baseline is trained with the same architecture, optimizer, number of updates, and ODE solver settings as the proposed method.
Third, beyond latent-space discrepancy and vector-field smoothness, we report image-space generative metrics after decoding. Specifically, we evaluate:
  • FID (lower is better), to assess overall sample quality with respect to the inlier-class reference set;
  • Recall (higher is better), to quantify diversity;
  • Coverage (higher is better), to assess mode support.
These metrics complement the latent-space RBF-MMD and the ambient roughness proxy
$\tilde{R}_{\mathrm{roughness}}(v_{\theta})=\mathbb{E}_{x\sim\mathcal{N}(0,I)}\bigl[\|\nabla_{x}v_{\theta}(x)\|_{F}^{2}\bigr],$
thereby allowing us to examine the trade-off between vector-field smoothness, robustness to off-manifold targets, and inlier-oriented decoded-image quality.
Fourth, we make the model-selection protocol explicit. We compare the following models under matched computational budgets:
$\mathrm{FM},\qquad\gamma\text{-}\mathrm{FM}\ (\text{shared-}t),\qquad\mathrm{FM+Jac},$
and optionally also report $\gamma$-FM++Jac when discussing whether density-weighting and explicit regularization are complementary. For fairness, the primary comparison is made under the same network, optimizer, training steps, and the same fixed-step midpoint solver budget. We select $\gamma$ from a fixed grid and $\lambda_{J}$ from a matched regularization grid using validation FID, and when possible we also report a matched-smoothness comparison to isolate the smoothness/diversity trade-off.
Table 1: Shared-time CIFAR-10 comparison under the robust-target latent protocol and matched computational budgets. All methods use the same latent autoencoder, the same flow backbone, the same number of training updates, and the same ODE solver budget.
[…]
FM (shared-$t$) 59.063 0.388 0.915 0.000609 0.012659 9.136
$\gamma$-FM ($\gamma=0.2$) 59.761 0.390 0.911 0.000631 0.012531 9.156
$\gamma$-FM ($\gamma=0.5$) 60.609 0.382 0.916 0.000685 0.012046 9.218
$\gamma$-FM ($\gamma=1.0$) 61.198 0.395 0.915 0.000773 0.011684 8.788
$\gamma$-FM ($\gamma=2.0$) 62.178 0.387 0.908 0.001060 0.011118 7.046
FM++Jac 60.305 0.397 0.924 0.000598 0.012915 9.709
Observed trade-off under the shared-time surrogate.
[…]
A complementary clean-target rerun helps interpret this result. In that setting, the weighting mechanism remained active, but the empirical advantage largely disappeared: shared-time FM was already competitive, and $\gamma$-FM became mostly comparable or slightly inferior in FID, Recall, and Coverage, while still tending to produce somewhat smoother vector fields. We therefore interpret the main gain of $\gamma$-weighting not as a universal improvement on all latent-generation tasks, but as a geometry-dependent benefit that becomes most visible when contaminated or low-support targets distort the regression problem.
[…]
Comparison with explicit Jacobian regularization.
The FM++Jac baseline is informative but exhibits a different profile from $\gamma$-FM. In the shared-time runs, explicit Jacobian penalization could improve in-support fit and in some cases slightly helped Recall or Coverage, but it did not reproduce the systematic reduction in out-of-support discrepancy obtained by density weighting. Its effect on smoothness and FID was also less consistent, and the seed-to-seed variability was generally larger. This suggests that Jacobian regularization acts mainly as a generic local smoother, whereas $\gamma$-weighting changes the regression geometry by selectively suppressing low-density updates. We therefore regard FM++Jac as a useful comparison baseline, but not as a replacement for the void-rejection mechanism induced by the density-weighted objective.
Adaptive selection via GSC and validation metrics.
To choose $\gamma$ without relying only on visual inspection, we continue to examine the Geometric Selection Criterion (GSC) from Section 3.6, but we interpret it as a practical heuristic rather than as a definitive optimality principle. In practice, we read the GSC curve together with validation FID and decoded-image diagnostics, so that the choice of $\gamma$ is supported by both latent-space and decoded-image summaries. In the present robust-target CIFAR-10 latent study, this combined view suggests $\gamma=0.5$ as a reasonable operating point, while $\gamma=1.0$ and $2.0$ are most useful for illustrating increasingly conservative low-density behavior.
Smoothness–diversity trade-off under the shared-time surrogate.
Figure 3 complements Table 1 by visualizing the same trade-off more directly. As $\gamma$ increases, the model becomes more conservative in low-density regions under the robust-target protocol, which improves the out-of-support discrepancy and the smoothness proxy, while increasing the in-support discrepancy. In this sense, the shared-time protocol reveals a coherent trade-off rather than a uniformly dominant improvement. Among the tested values, $\gamma=0.5$ provides a reasonable operating point in our robust-target latent CIFAR-10 experiments, whereas $\gamma=1.0$ and especially $\gamma=2.0$ make the same tendency more pronounced.
Viewed together with the clean-target sanity check, this figure also clarifies the regime in which the weighting is most useful. When the training targets are already clean, the same weighting mechanism still operates, but its effect is expressed mainly as additional regularization rather than as a clear gain in decoded-image quality. This supports the interpretation that the proposed method is particularly useful when the regression geometry is perturbed by contamination or by low-support target regions.
Figure 3: Shared-time $\gamma$-sweep on latent CIFAR-10 under the robust-target protocol and the rank-based piecewise surrogate. Left: $\mathrm{MMD}_{\mathrm{out}}^{2}$ decreases as $\gamma$ increases, indicating improved suppression of out-of-support trajectories. Middle: the smoothness proxy improves overall as $\gamma$ becomes larger, consistent with a more conservative vector field in low-density regions. Right: $\mathrm{MMD}_{\mathrm{in}}^{2}$ increases with $\gamma$, showing the accompanying cost in in-support fidelity. Together, these panels visualize the trade-off reported in Table 1. A separate clean-target rerun showed that the weighting mechanism remained active there as well, but the gain in decoded-image quality largely disappeared, so the present robust-target regime is the one in which the advantage of density weighting is most clearly expressed.